# Large-scale peculiar velocities in the universe

### v2

@@ 7.1.8 The generalised FRW case
since to zero order. The above holds on all FRW backgrounds, irrespective of the their spatial curvature and equation of state (with the exception of the models). Clearly, on subhorizon scales, where , the dominant (correction) term on the right-hand side of (7.1.28) is the one containing the aforementioned ratio. The rest of the terms have coefficients of order unity or less, given that and provided the density parameter () does not take unrealistically large values.232323Given that the perturbations , and are all very small during the linear regime, Eq. (7.1.28) suggests that on scales close and beyond the Hubble horizon. On these grounds, well inside the Hubble radius, the two deceleration parameters are still related by (7.1.13) and the transition length is still given by expression (7.1.14) of the Einstein-de Sitter background (see § 7.1.2 previously). Therefore, on subhorizon scales, the linear relative-motion effects on the deceleration parameter are the same on all physically realistic FRW backgrounds [249]. In other words, observers living inside (slightly) contracting bulk peculiar flows in any observationally realistic Friedmann universe are likely to misinterpret their slower local expansion rate as recent acceleration of the surrounding universe.
@@ 7.1.9 The Bianchi case
which reproduces (7.1.29a) once the expression is used for the peculiar flux. To linear order, the Bianchi background does not change the energy conservation law, which in the absence of pressure retains its almost-FRW form (see Eq. (7.1.8a) in § 7.1.1 previously). Taking the spatial gradient of (7.1.8a), using the linear commutation law and solving the resulting expression for the 4-acceleration yields [250]
This is the relativistic linear peculiar 4-acceleration in a tilted, almost-Bianchi universe filled with pressureless matter in the presence of peculiar velocities. Comparing the above to its almost-FRW counterpart (see Eq. (7.1.10) in § 7.1.1 previously), one notices that the Bianchi background has added a shear-induced term to the right-hand side.
@@
after a simple harmonic splitting. The impact of the Bianchi background is carried by the ratio, which also provides a measure of the model’s anisotropy. Assuming realistically anisotropic backgrounds we may set . Then, on subhorizon scales where , the two deceleration parameters are still related by Eq. (7.1.13) and the transition length is still given by expression (7.1.14) of the Einstein-de Sitter case (see § 7.1.2 previously).242424The Bianchi transition length, marking the threshold when , is generally given by [250] (7.1.35) and reduces to its Einstein de Sitter counterpart (7.1.14) when (as expected). There, as well as here, the transition length is a fraction of the Hubble scale and closely analogous to the Jeans length [81]. Here, the background anisotropy increases when and decreases it if (recall that ). In either case, the effect is weak in realistically anisotropic models (with ). Hence, as in the generalised FRW case (see § 7.1.8 before), on subhorizon scales the linear relative-motion effects on the deceleration parameter remain unaffected, unless the Bianchi background has unacceptably high anisotropy and/or unacceptably strong spatial curvature.252525In the observationally unrealistic case of high background anisotropy with , the relative-motion effects on are dominated by the aforementioned ratio. For example, confining to super-Hubble lengths, where , and allowing for high anisotropy, Eq. (7.1.34) reduces to (7.1.36) The effect is now solely determined by the sign of and by the magnitude of the correction term on the right-hand side of the above, which combines the kinematic anisotropy of the Bianchi background with that of the bulk peculiar flow. Then, when , the local value of the deceleration parameter decreases (i.e. ) and it can even drop below zero (see [250] for further discussion). Therefore, observers living inside (slightly) contracting bulk peculiar flows in observationally realistic Bianchi universes, are still likely to misinterpret their slower local expansion rate as recent acceleration of the surrounding universe.
@@ 7.1.10 The null case
where is the energy of the photons. The latter travel along null geodesics tangent to the -vector (with ) and they are received along the spacelike -direction by observers following timelike worldlines tangent to the -field.
@@
In an almost-FRW universe with peculiar-velocity perturbations, the null and the timelike deceleration parameters, as measured in the CMB and the tilted frames, are related by
@@
since in our FRW background. This means that the transition length for is approximately times larger than the one associated with its timelike counterpart () [253].
@@ 7.2.1 The deceleration tensor
given that . Hence, the deceleration parameter along a selected spatial direction consists of its mean value plus/minus shear-like corrections. As expected, in the spatially isotropic FRW universes, definitions (7.2.1) and (7.2.4) reduce to
@@ 7.4 Doppler-like dipoles in the number counts
Depending on their aims, typical cosmological studies employ two alternative reference frames. These are the energy (or Landau-Lifshitz) frame and the particle (or Eckart) frame, with associated 4-velocity vectors and respectively. In the former, the energy-flux vector is zero by default, while in the latter it is the particle-flux vector that vanishes (see decompositions (7.4.1) next). In a general spacetime the aforementioned coordinate systems are distinct, but in our homogeneous and isotropic FRW background, where all physical phenomena obey the (strict) cosmological principle, are are assumed to coincide.
Let us consider a distribution of astrophysical sources (e.g. supernovae, radio galaxies, etc) with 4-current density . Expressed relative to the CMB and the tilted frames, the 4-current density decomposes as
@@
Applying Eqs. (7.4.1) to a tilted almost-FRW universe and linearising, the two number densities and the corresponding flux vectors are related by
@@
Observationally speaking, the high isotropy of the CMB seems to provide irrefutable evidence that our universe is highly uniform on large enough scales. Following the theoretical arguments of [256], this belief has been recently put under question, after a number of reports claimed dipolar anisotropies in the number counts of distant sources that appeared in disagreement with their CMB counterpart, both in magnitude and in direction. It was the differences in the dipole amplitudes, in particular, that raised the major questions, which went as far as to doubt the validity of the Cosmological Principle itself. Such scenarios have also been proposed in the past, though primarily as theoretical alternatives.
The fact that large-scale relative motions can trigger (apparent) dipolar anisotropies in the number counts of distant astrophysical sources has been largely known. What is not so well known perhaps, is that the amplitudes of the aforementioned dipoles vary because they are proportional to the number densities of the associated sources. In addition, the directions of the induced dipolar axes differ from source to source, as well as from that of their CMB counterpart, due to evolution effects. All this should be taken into account when considering the inconsistencies reported in the number counts of distant astrophysical sources and their potential implications for the Cosmological Principle.
@@ 8.2.2 Multiple non-interacting perfect fluids
The differential formulae (8.2.12) and (8.2.14) monitor the linear evolution of a mixture of non-comoving, non-interacting barotropic perfect fluids on an spatially flat FRW background.303030For the system governing interacting, non-barotropic ideal media on a general Friedmann background, as well as for further discussion and additional references, the interested reader is referred to § 3.3.1, § 3.3.2 and § 3.3.3 of [88], or/and to § 10.3.2 of [89]. In order to close the system, one needs to specify the form of the peculiar fluxes ( and ). Next, we will do this within a perturbed spatially flat FRW background cosmology that contains a composition of radiation and pressure-free dust.
@@ 8.3 The Meszaros “stagnation” effect
given that to linear order and in the FRW background. The above immediately leads to
@@
This homogeneous differential equation is formally identical to the one governing the linear evolution of inhomogeneities in the dust on an Einstein-de Sitter background. Here, Eq. (8.3.4) holds on a spatially flat FRW background filled with a non-interacting radiation-dust mixture and monitors the linear evolution of inhomogeneities in the dust component, assuming that the radiation field is homogeneously distributed in space. We may therefore use (8.3.4) to infer the linear evolution of both after and prior to equipartition.
@@
Both (8.3.5) and its associated solution (8.3.6) are identical to their Newtonian counterparts (e.g. see [87, 260] for the respective expressions). However, this agreement could misleading, since the relativistic analysis presented here degenerated (for all practical purposes) to Newtonian once the Landau-Lifshitz frame was adopted and the gravitational input of the peculiar flux was bypassed. In view of this, it would be interesting to see whether the Newtonian picture reproduced in Eqs. (8.3.5) and (8.3.6) changes when the aforementioned flux-input is properly included in the relativistic treatment of multi-component media.343434In cosmological perturbation theory two are the typical reference coordinate systems: the Landau-Lifshitz frame and the Eckart frame [88, 89]. The former sets the energy flux to zero, while the latter switches the particle flux off. This means that adopting the Landau-Lifshitz frame when studying a perturbed multi-component medium may set the energy flux to zero, but triggers a nonzero effective particle flux. Using the Eckart frame, on the other hand does the exact opposite. In either case, a system of several non-comoving perfect fluids cannot be treated as perfect.
With the comments of the last paragraph in mind, we will re-iterate to the reader our earlier statement (see § 8.1 previously) that, within the Newtonian/quasi-Newtonian framework, peculiar velocities leave the linear evolution of the (post-recombination) density gradients unaffected.
@@ 9.5.2 The H-dipole
In a follow up study [58], the authors analysed an increased sample of up to 570 clusters with measured X-ray, infrared and microwave properties. This significantly improved the sky-coverage and the statistical strength of the study. The new data were used to construct ten different scaling relations, in addition to , in order to test the isotropy of the Hubble expansion. Employing the same methodology with [57], the results suggested a 9% variation in the value of the Hubble constant between ) and the rest of the sky (see Fig. 19 above). The reported anisotropy had a nearly dipole form and statistical significance greater than [58]. Although the authors could not entirely exclude interference from as yet unknown systematics, they felt confident to claim that their results do not suffer from any known biases. After all, it would be quite unlikely for the systematics to bias all the scaling relations simultaneously and lead to the same results across different observables and samples. Without excluding the possibility the reported -dipole to reflect a generic large-scale (cosmic) anisotropy of the universe, the authors suggested that the observed asymmetry could be due to a bulk flow extending out to more than 500 Mpc and moving with 800 km/sec (much like the dark flows reported by Kashlinsky et al - see § 9.4 previously). Both alternatives are in serious tension with the standard CDM model.
@@ 9.5.3 Number-count dipoles
The Cosmological Principle (CP), namely the belief that we are all typical observers in a nearly homogeneous and isotropic (i.e. almost-FRW) universe has been the cornerstone of modern cosmology for the last hundred years. The strongest observational support for the CP so far, comes form the CMB spectrum, which exhibits a remarkably high degree of isotropy. Combined with the Copernican Principle, advocating that we are not privileged observers in the cosmos, the isotropy of the CMB leads to the conclusion that we live in a maximally (i.e. spatially homogeneous and isotropic) universe. There is also theoretical support for CP, primarily coming from the inflationary paradigm, which is supposed to dilute essentially any degree of primordial anisotropy and/or inhomogeneity. Although structure formation has destroyed (to a lesser or larger degree) the overall uniformity of the universe on relatively small scales, the established belief maintains that beyond a certain length our cosmos remains largely homogeneous and isotropic. The uniformity threshold, however, which is typically set close to the 100 Mpc mark, is uncertain and under debate.
@@ 10.1.1 Redshift and luminosity distance measurements
Let us now consider a tilted observer, with 4-velocity and peculiar velocity . Assuming that the peculiar motion is non-relativistic, namely that , the above satisfy the linear Lorentz boost (6.3.1). Applying the latter to the tilted frame we may write , which combines with (10.1.1) to give
since the photon energy at emission is the same in both frames (i.e. ). To proceed further, one projects (6.3.1) along the null direction to obtain , which immediately translates into . Here, and represent the photon energies measured in the reference (CMB) frame and in the tilted (matter) frames respectively. Using (7.1.37) and keeping up to -order terms, the relation between the two photon energies reads
with being the unit spacelike vector along the direction of the emission (see § 7.1.10 previously). Therefore, the scalar measures the change in the photon energy triggered by the 4-velocity boost (6.3.1), that is by the tilted observer’s peculiar flow. Finally, going back to relation , we immediately obtain
@@
to first approximation (recall that the index 2 marks the time of reception). The above monitors the effect of weak peculiar motions on the measured redshift of a distant luminous source.
Effects on the luminosity distance: Perhaps the most familiar impact of peculiar velocities is the one on distance measurements. For instance, following [310], in an FRW cosmology with no peculiar velocities the luminosity distance of a source at redshift is given by
@@
Allowing for (weak) peculiar velocities, but keeping the FRW nature of the host spacetime unchanged, recasts (10.1.6) into [310]
@@
after dropping terms of order higher than and . The above expression provides the luminosity distance of a source located at redshift in an FRW universe endowed with a weak peculiar-velocity field. Following () and (10.1.11) and (10.1.12), the effect of teh peculiar motion on the luminosity distance vanishes when , or equivalently when (see related discussion above). Alternatively, one may say that there is no relative-motion effect on the luminosity distance when the peculiar velocity is normal to the line-of-sight (see also Eq. (10.1.14) below).

### v3

@@ 7.1.8 The generalised FLRW case
since to zero order. The above holds on all FLRW backgrounds, irrespective of the their spatial curvature and equation of state (with the exception of the models). Clearly, on subhorizon scales, where , the dominant (correction) term on the right-hand side of (7.1.28) is the one containing the aforementioned ratio. The rest of the terms have coefficients of order unity or less, given that and provided the density parameter () does not take unrealistically large values.232323Given that the perturbations , and are all very small during the linear regime, Eq. (7.1.28) suggests that on scales close and beyond the Hubble horizon. On these grounds, well inside the Hubble radius, the two deceleration parameters are still related by (7.1.13) and the transition length is still given by expression (7.1.14) of the Einstein-de Sitter background (see § 7.1.2 previously). Therefore, on subhorizon scales, the linear relative-motion effects on the deceleration parameter are the same on all physically realistic FLRW backgrounds [272]. In other words, observers living inside (slightly) contracting bulk peculiar flows in any observationally realistic Friedmann universe are likely to misinterpret their slower local expansion rate as recent acceleration of the surrounding universe.
@@ 7.1.9 The Bianchi case
which reproduces (7.1.29a) once the expression is used for the peculiar flux. To linear order, the Bianchi background does not change the energy conservation law, which in the absence of pressure retains its almost-FLRW form (see Eq. (7.1.8a) in § 7.1.1 previously). Taking the spatial gradient of (7.1.8a), using the linear commutation law and solving the resulting expression for the 4-acceleration yields [273]
This is the relativistic linear peculiar 4-acceleration in a tilted, almost-Bianchi universe filled with pressureless matter in the presence of peculiar velocities. Comparing the above to its almost-FLRW counterpart (see Eq. (7.1.10) in § 7.1.1 previously), one notices that the Bianchi background has added a shear-induced term to the right-hand side.
@@
after a simple harmonic splitting. The impact of the Bianchi background is carried by the ratio, which also provides a measure of the model’s anisotropy. Assuming realistically anisotropic backgrounds we may set . Then, on subhorizon scales where , the two deceleration parameters are still related by Eq. (7.1.13) and the transition length is still given by expression (7.1.14) of the Einstein-de Sitter case (see § 7.1.2 previously).242424The Bianchi transition length, marking the threshold when , is generally given by [273] (7.1.35) and reduces to its Einstein de Sitter counterpart (7.1.14) when (as expected). There, as well as here, the transition length is a fraction of the Hubble scale and closely analogous to the Jeans length [92]. Here, the background anisotropy increases when and decreases it if (recall that ). In either case, the effect is weak in realistically anisotropic models (with ). Hence, as in the generalised FLRW case (see § 7.1.8 before), on subhorizon scales the linear relative-motion effects on the deceleration parameter remain unaffected, unless the Bianchi background has unacceptably high anisotropy and/or unacceptably strong spatial curvature.252525In the observationally unrealistic case of high background anisotropy with , the relative-motion effects on are dominated by the aforementioned ratio. For example, confining to super-Hubble lengths, where , and allowing for high anisotropy, Eq. (7.1.34) reduces to (7.1.36) The effect is now solely determined by the sign of and by the magnitude of the correction term on the right-hand side of the above, which combines the kinematic anisotropy of the Bianchi background with that of the bulk peculiar flow. Then, when , the local value of the deceleration parameter decreases (i.e. ) and it can even drop below zero (see [273] for further discussion). Therefore, observers living inside (slightly) contracting bulk peculiar flows in observationally realistic Bianchi universes, are still likely to misinterpret their slower local expansion rate as recent acceleration of the surrounding universe.
@@ 7.1.10 The null case
where is the energy of the photons. The latter travel along null geodesics tangent to the -vector (with ) and they are received along the spacelike -direction by observers following timelike worldlines tangent to the -field (e.g. see Fig. 7).
In what follows, we will assume that the radiation source is sufficiently distant, namely it has sufficiently high redshift, to be treated as comoving with the universal expansion. We may then ignore the peculiar velocity of the source at the point of emission and align its 4-velocity there with that of the CMB frame (). The signal is received by the realistic (tilted) observer (red worldline) at the same future event, where it is causally boosted to the CMB frame (blue worldline). Then, any difference in the measurements (e.g. like that seen in Eq. (7.1.39) below) will be attributed to the peculiar motion of the realistic observer (see also Fig. 7).
@@
In an almost-FLRW universe with peculiar-velocity perturbations, the null and the timelike deceleration parameters, as measured in the CMB and the tilted frames simultaneously at the reception event (see Fig. 7), are related by
@@
since in our FLRW background. This means that the transition length for is approximately times larger than the one associated with its timelike counterpart () [276].
@@ 7.2.1 The deceleration tensor
given that . Hence, the deceleration parameter along a selected spatial direction consists of its mean value plus/minus shear-like corrections. As expected, in the spatially isotropic FLRW universes, definitions (7.2.1) and (7.2.4) reduce to
@@ 7.4 Doppler-like dipoles in the number counts
Observationally speaking, the high isotropy of the CMB seems to provide irrefutable evidence that our universe is highly uniform on large enough scales. Following the theoretical arguments of [278], however, this belief has been recently put under question. The reason can be traced to number of reports claimed dipolar anisotropies in the number counts of distant sources that appeared in disagreement with their CMB counterpart both in magnitude and in direction. It was the differences in the dipole amplitudes, in particular, that raised the major questions, which went as far as to doubt the validity of the Cosmological Principle itself. Such scenarios have also been proposed in the past, though primarily as theoretical alternatives. The reader is referred to § 9.5.3 here for a brief discussion of the observations and to the recent review of [263] for details and more references.
The question of the Cosmological Principle lies outside the scope of this review. We will therefore outline here the results of the simple, though informative, study of [279] on the peculiar-velocity effects on the number counts of distant astrophysical sources, referring the reader to studies, such as those of [280, 281] for example, for more extensive treatments.
Depending on their aims, typical cosmological studies employ two alternative reference frames. These are the energy (or Landau-Lifshitz) frame and the particle (or Eckart) frame, with associated 4-velocity vectors and respectively. In the former, the energy-flux vector is zero by default, while in the latter it is the particle-flux vector that vanishes (see decompositions (7.4.1) next). In a general spacetime the aforementioned coordinate systems are distinct, but in our homogeneous and isotropic FLRW background, where all physical phenomena obey the (strict) cosmological principle, are are assumed to coincide.
Let us consider a distribution of astrophysical sources (e.g. supernovae, radio galaxies, etc) with 4-current density . As in § 7.1.10 before, the sources are at sufficiently high redshifts to ignore their individual peculiar velocities and treat them as comoving. Also, the number counts are measured locally (i.e. at the same spacetime event) by the tilted and the CMB observers. Expressed relative to the CMB and the tilted frames, the 4-current density decomposes as
@@
Applying Eqs. (7.4.1) to a tilted almost-FLRW universe and linearising, the two number densities and the corresponding flux vectors are related by
@@
The fact that large-scale relative motions can trigger (apparent) dipolar anisotropies in the number counts of distant astrophysical sources has been largely known (e.g. see [279]). What is not so well known perhaps, is that the amplitudes of the aforementioned dipoles vary because they are proportional to the number densities of the associated sources. Such differences may come to lie within our observational capabilities soon, if not already. In addition, the directions of the induced dipolar axes differ from source to source, as well as from that of their CMB counterpart, due to evolution effects. All this should be taken into account when considering the inconsistencies reported in the number counts of distant astrophysical sources and their potential implications for the Cosmological Principle.
@@ 8.2.2 Multiple non-interacting perfect fluids
The differential formulae (8.2.12) and (8.2.14) monitor the linear evolution of a mixture of non-comoving, non-interacting barotropic perfect fluids on an spatially flat FLRW background.303030For the system governing interacting, non-barotropic ideal media on a general Friedmann background, as well as for further discussion and additional references, the interested reader is referred to § 3.3.1, § 3.3.2 and § 3.3.3 of [99], or/and to § 10.3.2 of [100]. In order to close the system, one needs to specify the form of the peculiar fluxes ( and ). Next, we will do this within a perturbed spatially flat FLRW background cosmology that contains a composition of radiation and pressure-free dust.
@@ 8.3 The Meszaros “stagnation” effect
given that to linear order and in the FLRW background. The above immediately leads to
@@
This homogeneous differential equation is formally identical to the one governing the linear evolution of inhomogeneities in the dust on an Einstein-de Sitter background. Here, Eq. (8.3.4) holds on a spatially flat FLRW background filled with a non-interacting radiation-dust mixture and monitors the linear evolution of inhomogeneities in the dust component, assuming that the radiation field is homogeneously distributed in space. We may therefore use (8.3.4) to infer the linear evolution of both after and prior to equipartition.
@@
Both (8.3.5) and its associated solution (8.3.6) are identical to their Newtonian counterparts (e.g. see [98, 285] for the respective expressions). However, this agreement could misleading, since the relativistic analysis presented here degenerated (for all practical purposes) to Newtonian once the Landau-Lifshitz frame was adopted and the gravitational input of the peculiar flux was bypassed.343434In cosmological perturbation theory two are the typical reference coordinate systems: the Landau-Lifshitz frame and the Eckart frame [99, 100]. The former sets the energy flux to zero, while the latter switches the particle flux off. This means that adopting the Landau-Lifshitz frame when studying a perturbed multi-component medium may set the energy flux to zero, but triggers a nonzero effective particle flux. Using the Eckart frame, on the other hand does the exact opposite. In either case, a system of several non-comoving perfect fluids cannot be treated as perfect. In view of this, it would be interesting to see whether the Newtonian picture reproduced in Eqs. (8.3.5) and (8.3.6) changes when the aforementioned flux-input is properly included in the relativistic studies of the peculiar-velocity effects on the evolution of density perturbations.
Although such a relativistic analysis is still missing, a recent essay has raised the possibility that structure formation could proceed more efficiently when the effects of peculiar motions are accounted for [286]. This could in turn lead to fully formed galaxies at redshifts higher than it is generally anticipated. Perhaps like those reported on the basis of the James Webb Space Telescope (JWST) data [287, 288, 289, 290, 291], which seems at odds with the CDM model [292, 293, 294].
@@ 9.5.2 The H-dipole
In a follow up study [65], the authors analysed an increased sample of up to 570 clusters with measured X-ray, infrared and microwave properties. This significantly improved the sky-coverage and the statistical strength of the study. The new data were used to construct ten different scaling relations, in addition to , in order to test the isotropy of the Hubble expansion. Employing the same methodology with [64], the results suggested a 9% variation in the value of the Hubble constant between ) and the rest of the sky (see Fig. 20 above). The reported anisotropy had a nearly dipole form and statistical significance greater than [65]. Although the authors could not entirely exclude interference from as yet unknown systematics, they felt confident to claim that their results do not suffer from any known biases. After all, it would be quite unlikely for the systematics to bias all the scaling relations simultaneously and lead to the same results across different observables and samples. Without excluding the possibility the reported -dipole to reflect a generic large-scale (cosmic) anisotropy of the universe (a possibility raised in [160, 338]), the authors suggested that the observed asymmetry could be due to a bulk flow extending out to more than 500 Mpc and moving with 800 km/sec (much like the dark flows reported by Kashlinsky et al - see § 9.4 previously). Both alternatives are in serious tension with the standard CDM model.
@@ 9.5.3 Number-count dipoles
The Cosmological Principle (CP), namely the belief that we are all typical observers in a nearly homogeneous and isotropic (i.e. almost-FLRW) universe has been the cornerstone of modern cosmology for the last hundred years. The strongest observational support for the CP so far, comes form the CMB spectrum, which exhibits a remarkably high degree of isotropy. Combined with the Copernican Principle, advocating that we are not privileged observers in the cosmos, the isotropy of the CMB leads to the conclusion that we live in a maximally (i.e. spatially homogeneous and isotropic) universe. There is also theoretical support for CP, primarily coming from the inflationary paradigm, which is supposed to dilute essentially any degree of primordial anisotropy and/or inhomogeneity. Although structure formation has destroyed (to a lesser or larger degree) the overall uniformity of the universe on relatively small scales, the established belief maintains that beyond a certain length our cosmos remains largely homogeneous and isotropic. The uniformity threshold, however, which is typically set close to the 100 Mpc mark, is uncertain and under debate.
@@ 10.1.1 Redshift and luminosity distance measurements
Let us now consider a pair of idealised (CMB) and realistic (tilted) observers, with respective 4-velocities and , located at the reception point. Relative to the tilted frame Eq. (10.1.1) reads , since the photon energy at the emission point () is unchanged. The latter combines with (10.1.1) to give
Note that, as stated in § 7.1.10 and § 7.4 before, the involved sources are assumed to have sufficiently high redshifts to ignore their individual peculiar velocities at the point of emission and treat them as comoving with the CMB frame and the universal expansion.
To proceed further, one projects (6.3.1) along the null direction to obtain , which immediately translates into . Here, and represent the photon energies measured in the reference (CMB) frame and in the tilted (matter) frames respectively. Using (7.1.37) and keeping up to -order terms, the relation between the two photon energies reads
with being the unit spacelike vector in the direction of the source (see § 7.1.10 previously). Therefore, the scalar measures the change in the photon energy triggered by the 4-velocity boost (6.3.1), that is by the tilted observer’s peculiar flow. Finally, going back to relation , we immediately obtain
@@
to first approximation (recall that the index 2 marks the time of reception). In the above and are the redshifts of the distant luminous source, as measured by the CMB and the tilted observer at the same spacetime event. Therefore, expression (10.1.5) provides the familiar redshift correction due to the observer’s peculiar motion relative to the CMB (e.g. see Eqs. (5), (6) in [345] for comparison). Note that, in the literature, is also referred to as the total, or the observed, or the heliocentric redshift, while is also known as the cosmological, or the true redshift. Finally, the scalar is occasionally referred to as the peculiar, or the Doppler, redshift.
Note that here we have isolated and focused upon the effect of the tilted observer’s peculiar motion. The redshift of a distant source is also distorted by additional effects, that occur during the propagation of the radiation signal, like the gravitational redshift and the integrated Sachs-Wolfe effects for example (e.g. see [346]).
Effects on the luminosity distance: An additional familiar impact of peculiar velocities is the one on distance measurements. For instance, following [347], in an FLRW cosmology with no peculiar velocities the luminosity distance of a source at redshift is given by
@@
Allowing for (weak) peculiar velocities, but keeping the FLRW nature of the host spacetime unchanged, recasts (10.1.6) into [347]
@@
after dropping terms of order higher than and . The above expression provides the luminosity distance of a source located at redshift in an FLRW universe endowed with a weak peculiar-velocity field. Following () and (10.1.11) and (10.1.12), the effect of teh peculiar motion on the luminosity distance vanishes when , or equivalently when (see related discussion above). Alternatively, one may say that there is no relative-motion effect on the luminosity distance when the peculiar velocity is normal to the line-of-sight (see also Eq. (10.1.14) below).